Hello,
I am stumped on this one please help,
If the area of a circle is 72 cm^2, find the radius to the nearest
hundredth.
Thanks!

Answer by theidiot(3)

the formula for the area of a square is A = pi * r^2
72 = pi * r^2
pi = 3.14 or 22/7
or use a calcuator and press the pi button
72 = pi * r^2
22.9183 = r^2
4.7873 = r
r = 4.79 cm